# On the smoothability problem with rational coefficients

### v1

@@ 0.1. Smoothing cycles
The strongest result in that direction, due to Kollár and Voisin [KV23, Theorem 1.2] (following an earlier work of Hironaka [Hir68, Theorem, Section 5, p. 50]), states that the answer to this question is positive in the range $c>\frac{n}{2}$. Additional positive results have been obtained by Kleiman [Kle69, Theorem 5.8] when $c=2$ and $n\in\{4,5\}$.
It was however discovered by Hartshorne, Rees and Thomas [HRT74, Theorem 1] that the answer to the question of Borel and Haefliger is negative in general. Further counterexamples have appeared in [Deb95, Théorème 6], in [Ben20, Theorem 0.3] (which includes examples in the threshold case $c=\frac{n}{2}$), and in [BD23, Theorem 1.2].
@@ 0.2. Rational coefficients
The above question was investigated by Kleiman in [Kle69] by exploiting the formula ${c_{c}(\mathcal{O}_{Z})=(-1)^{c}(c-1)!\,[Z]}$, a well-known consequence of the Grothendieck–Riemann–Roch theorem valid for all codimension $c$ subvarieties $Z\subset X$. Using this formula, one can reduce Question 0.1 to the case of Chern classes of vector bundles, which come by pull-back from Grassmannians (see the discussion in [Kle69, p. 282]). On the one hand, this reduces Question 0.1 to the case where $X$ is a Grassmannian. On the other hand, combined with an analysis of the singularities of the Schubert subvarieties of Grassmannians, this line of reasoning allowed Kleiman to show that Question 0.1 has a positive answer when $c>\frac{n}{2}-1$ (see [Kle69, Theorem 5.8]).
@@
0.3. Relation with Harthorne’s conjecture
@@ 0.4. A symplectic analogue
The proof of Theorem 0.4 is given in Section 2. We construct topological complex vector bundles on $M$ by means of Sullivan’s rational homotopy theory (in §2.1), and we apply the Auroux–Donaldson theorem to find sections of these bundles whose zero loci are symplectic submanifolds (in §2.2).
@@ 1.1. Statement
Since $n_{0}\geq 5$, one has $n\geq 10$. A Barth–Lefschetz-type theorem due to Sommese (see [Som82, Proposition 3.4 and (3.6.3)]) therefore shows that if $Z$ as in the corollary, then the restriction maps $H^{l}(G(k,n),{\mathbb{Z}})\to H^{l}(Z,{\mathbb{Z}})$ are bijective for $l\leq 2$, and hence that $\mathrm{Pic}(Z)$ generated by $\mathcal{L}_{\mid Z}$. In particular $Z$ is subcanonical and, as $H^{2}(G(k,n),\mathcal{L}^{i})=0$ for all $i$, the Serre construction applies, providing a rank 2 vector bundle $E$ such that $Z$ is the zero-locus of a section of $E$. The class of $Z$ is then equal to $c_{2}(E)$ so the corollary follows from Proposition 1.2. ∎
@@
The pair $\{a_{W},b_{W}\}$ is in fact independent of $W$.
@@
given by the universal family of $\mathbb{P}^{k}_{W}$’s, where $p$ is a projective bundle with fiber $\mathbb{P}^{k}_{W}$ over the point $[W]\in G(k+1,n)$. We denote by $\mathcal{L}$ and $\mathcal{L}^{\prime}$ the Plücker line bundles on $G(k,n)$ and $G(k+1,n)$. As we have $h^{0}(E_{\mid\mathbb{P}^{k}_{W}}(-a))=1$ for any $[W]\in G(k+1,n)$, the sheaf $R^{0}p_{*}(q^{*}(E\otimes\mathcal{L}^{-a}))$ is a line bundle on $G(k+1,n)$, hence isomorphic to ${\mathcal{L}^{\prime}}^{\otimes l}$ for some integer $l$. Furthermore, the natural inclusion
@@ 1.3. The unstable variant
In this statement, “unstable" means “not slope stable", and this condition says that there exists a nonzero section of $E(-a)$ for some $a$ such that ${\rm det}\,E(-a)\leq 0$. Conjecture 1.6 is important and, as explained in [Sch80], it would have remarkable consequences, such as the existence of topological complex vector bundles (of rank $2$) on projective space that do not have an algebraic structure. This conjecture is studied in [GS77] and [Siu15] but the proofs there have serious gaps, hence it is still open.
@@ 1.4. Extension to higher rank bundles
If Question 1.9 has a positive answer for $r\geq 2$, then for any rank $r$ vector bundle $E$ on $G(k,n)$ with $n-k\geq k\geq n_{0}(r)$, the Chern classes $c_{i}(E)$ for $i=1,\ldots,r$ satisfy
(1.11) $\int_{G(k,n)}c_{i}(E)l^{n-k-i}(l^{e}\gamma_{k}-\gamma_{n-k})=0,$
where $e=n-2k$, $l=c_{1}(\mathcal{L})$ and
@@
a positive answer to Question 1.9 implies that the restrictions $E_{\mid\mathbb{P}^{k}_{W}}$ and $E_{\mid\mathbb{P}^{n-k}_{W^{\prime}}}$ are decomposable vector bundles and by the same arguments as in the proof of Lemma 1.4, we conclude that the type of the decomposition is the same for $E_{\mid\mathbb{P}^{k}_{W}}$ and $E_{\mid\mathbb{P}^{n-k}_{W^{\prime}}}$ (and in particular it does not depend on the choice of $W$ and $W^{\prime}$). Hence there exist integers $a_{1},\ldots,\,a_{r}$ such that
@@
(1.14) $\int_{G(k,n)}l^{k-i}c_{i}(E)\gamma_{k}=\int_{\mathbb{P}^{k}_{W}}l^{k-i}c_{i}(E)=\sigma_{i}(a_{j})$
$\int_{G(k,n)}l^{n-k-i}c_{i}(E)\gamma_{n-k}=\int_{\mathbb{P}^{n-k}_{W^{\prime}}}l^{k-i}c_{i}(E)=\sigma_{i}(a_{j}),$
@@
The following lemma shows that the equations (1.11) impose $r-1$ independent equations to the Chern classes of $E$ which are thus severely restricted if Question 1.9 has a positive answer.
@@
of cup-product by the class $l^{n-k-i}(l^{e}\gamma_{k}-\gamma_{n-k})$ is nonzero for $i\geq 2$. Equivalently, thanks to Poincaré duality, the class $l^{n-k-i}(l^{e}\gamma_{k}-\gamma_{n-k})$ is nonzero for $i\geq 2$.
@@
It clearly suffices to prove that the cohomology class $l^{n-k-2}(l^{e}\gamma_{k}-\gamma_{n-k})$ is nonzero in $H^{2N-4}(G(k,n),\mathbb{Z})$. This class is the difference of the class of a plane in $\mathbb{P}^{k}_{W}$ and the class of a plane in $\mathbb{P}^{n-k}_{W^{\prime}}$. We construct an inclusion
@@ 2.2. The Auroux–Donaldson theorem
To be precise, the Auroux–Donaldson formulation assumes that the de Rham cohomology class $\frac{[\omega]}{2\pi}$ is is in the image of the natural map $H^{2}(M,{\mathbb{Z}})\to H^{2}(M,{\mathbb{R}})$. This restriction is well-known to be immaterial, as indicated in [Don96, Corollary 6] and [Sik98, Remarque p. 233]. For the convenience of the reader, we give a few more details on this argument.

### v2

@@ 0.1. Smoothing cycles
The strongest result in that direction, due to Kollár and Voisin [KV24, Theorem 1.2] (following an earlier work of Hironaka [Hir68, Theorem, Section 5, p. 50]), states that the answer to this question is positive in the range $c>\frac{n}{2}$. Additional positive results have been obtained by Kleiman [Kle69, Theorem 5.8] when $c=2$ and $n\in\{4,5\}$.
It was however discovered by Hartshorne, Rees and Thomas [HRT74, Theorem 1] that the answer to the question of Borel and Haefliger is negative in general. Further counterexamples have appeared in [Deb95, Théorème 6], in [Ben24, Theorem 0.3] (which includes examples in the threshold case $c=\frac{n}{2}$), and in [BD23, Theorem 1.2].
@@ 0.2. Rational coefficients
The above question was investigated by Kleiman in [Kle69] by exploiting the formula ${c_{c}(\mathcal{O}_{Z})=(-1)^{c}(c-1)!\,[Z]}$ for the $c$-th Chern class of the structure sheaf of a codimension $c$ subvariety $Z\subset X$ (a well-known consequence of the Grothendieck–Riemann–Roch theorem). Using this formula, one can reduce Question 0.1 to the case of Chern classes of vector bundles, which come by pull-back from Grassmannians (see the discussion in [Kle69, p. 282]). On the one hand, this reduces Question 0.1 to the case where $X$ is a Grassmannian. On the other hand, combined with an analysis of the singularities of the Schubert subvarieties of Grassmannians, this line of reasoning allowed Kleiman to show that Question 0.1 has a positive answer when $c>\frac{n}{2}-1$ (see [Kle69, Theorem 5.8]).
@@
0.3. Relation with Hartshorne’s conjecture
@@ 0.4. A symplectic analogue
The proof of Theorem 0.4 is given in Section 2. We construct in §2.1 topological complex vector bundles on $M$ by means of Sullivan’s rational homotopy theory, and we apply in §2.2 the Auroux–Donaldson theorem to find sections of these bundles whose zero loci are symplectic submanifolds.
@@ 1.1. Statement
Since $n_{0}\geq 5$, one has $n\geq 10$. A Barth–Lefschetz-type theorem due to Sommese (see [Som82, Proposition 3.4 and (3.6.3)]) therefore shows that if $Z$ is as in the corollary, then the restriction maps $H^{l}(G(k,n),{\mathbb{Z}})\to H^{l}(Z,{\mathbb{Z}})$ are bijective for $l\leq 2$, and hence that $\mathrm{Pic}(Z)$ is generated by $\mathcal{L}_{\mid Z}$. In particular $Z$ is subcanonical and, as $H^{2}(G(k,n),\mathcal{L}^{i})=0$ for all $i$, the Serre construction applies, providing a rank 2 vector bundle $E$ such that $Z$ is the zero-locus of a section of $E$. The class of $Z$ is then equal to $c_{2}(E)$ so the corollary follows from Proposition 1.2. ∎
@@
The pair $\{a_{W},b_{W}\}$ is in fact independent of $W$ and we have $\{a_{W},b_{W}\}=\{a_{W^{\prime}},b_{W^{\prime}}\}$.
@@
Remark 1.6.
The proof of Lemma 1.4 shows more generally that any rank $r$ vector bundle on $G(k,n)$ that is split on each $\mathbb{P}^{k}_{W}$ and $\mathbb{P}^{n-k}_{W^{\prime}}$ has constant splitting type and the splitting type is the same for $\mathbb{P}^{k}_{W}$ and $\mathbb{P}^{n-k}_{W^{\prime}}$. If $r\leq k$, there is an even easier proof of the first statement. Namely, the constancy follows from the fact that, for a split vector bundle of rank $r$ on $\mathbb{P}^{k}$ with $r\leq k$, its splitting type is determined by its Chern classes.
@@
given by the universal family of $\mathbb{P}^{k}_{W}$’s, where $p$ is a projective bundle with fiber $\mathbb{P}^{k}_{W}$ over the point $[W]\in G(k+1,n)$. We denote respectively by $\mathcal{L}$ and $\mathcal{L}^{\prime}$ the Plücker line bundles on $G(k,n)$ and $G(k+1,n)$. As we have $h^{0}(E_{\mid\mathbb{P}^{k}_{W}}(-a))=1$ for any $[W]\in G(k+1,n)$, the sheaf $R^{0}p_{*}(q^{*}(E\otimes\mathcal{L}^{-a}))$ is a line bundle on $G(k+1,n)$, hence isomorphic to ${\mathcal{L}^{\prime}}^{\otimes l}$ for some integer $l$. Furthermore, the natural inclusion
@@ 1.3. The unstable variant
In this statement, “unstable” means “not slope stable”, and this condition says that there exists a nonzero section of $E(-a)$ for some $a$ such that ${\rm det}\,E(-a)\leq 0$. Conjecture 1.7 is important and, as explained in [Sch80], it would have remarkable consequences, such as the existence of topological complex vector bundles (of rank $2$) on projective space that do not have an algebraic structure. This conjecture is studied in [GS77] and [Siu15] but the proofs there have serious gaps, hence it is still open.
@@ 1.4. Extension to higher rank bundles
If Question 1.10 has a positive answer for $r\geq 2$, then for any rank $r$ vector bundle $E$ on $G(k,n)$ with $n-k\geq k\geq n_{0}(r)$, the Chern classes $c_{i}(E)$ for $i=1,\ldots,r$ satisfy
(1.11) $\int_{G(k,n)}c_{i}(E)l^{k-i}(\gamma_{k}-l^{n-2k}\gamma_{n-k})=0,$
where $l=c_{1}(\mathcal{L})$ and
@@
a positive answer to Question 1.10 implies that the restrictions $E_{\mid\mathbb{P}^{k}_{W}}$ and $E_{\mid\mathbb{P}^{n-k}_{W^{\prime}}}$ are decomposable vector bundles and by the same arguments as in the proof of Lemma 1.4 (see Remark 1.6), we conclude that the type of the decomposition is the same for $E_{\mid\mathbb{P}^{k}_{W}}$ and $E_{\mid\mathbb{P}^{n-k}_{W^{\prime}}}$ (and in particular it does not depend on the choice of $W$ and $W^{\prime}$). Hence there exist integers $a_{1},\ldots,\,a_{r}$ such that
@@
(1.14) $\int_{G(k,n)}l^{k-i}c_{i}(E)\gamma_{k}=\int_{\mathbb{P}^{k}_{W}}h^{k-i}c_{i}(E)=\sigma_{i}(a_{j})$
$\int_{G(k,n)}l^{n-k-i}c_{i}(E)\gamma_{n-k}=\int_{\mathbb{P}^{n-k}_{W^{\prime}}}(h^{\prime})^{n-k-i}c_{i}(E)=\sigma_{i}(a_{j}),$
@@
The following lemma shows that the equations (1.11) impose $r-1$ independent equations to the Chern classes of $E$ which are thus severely restricted if Question 1.10 has a positive answer.
@@
of cup-product by the class $l^{k-i}(\gamma_{k}-l^{n-2k}\gamma_{n-k})$ is nonzero for $i\geq 2$. Equivalently, thanks to Poincaré duality, the class $l^{k-i}(\gamma_{k}-l^{n-2k}\gamma_{n-k})$ is nonzero for $i\geq 2$.
@@
It clearly suffices to prove that the cohomology class $l^{k-2}(\gamma_{k}-l^{n-2k}\gamma_{n-k})$ is nonzero in $H^{2N-4}(G(k,n),\mathbb{Z})$. This class is the difference of the class of a plane in $\mathbb{P}^{k}_{W}$ and the class of a plane in $\mathbb{P}^{n-k}_{W^{\prime}}$. We construct an inclusion
@@ 2.2. The Auroux–Donaldson theorem
To be precise, the Auroux–Donaldson formulation assumes that the de Rham cohomology class $\frac{[\omega]}{2\pi}$ is in the image of the natural map $H^{2}(M,{\mathbb{Z}})\to H^{2}(M,{\mathbb{R}})$. This restriction is well-known to be immaterial, as indicated in [Don96, Corollary 6] and [Sik98, Remarque p. 233]. For the convenience of the reader, we give a few more details on this argument.